# SPECIAL RELATIVITY

### v11

@@ 8.3 Measuring Wigner rotation
Here $d\vec{v}$ represents the infinitesimal change in velocity due to acceleration, $d\Phi$ is the Wigner rotation angle describing the rotation of the particle’s proper frame relative to the Lorentz lab frame, and $d\theta$ is the infinitesimal orbital angle. A natural question arises: how can this rotation be observed or measured? Since a moving coordinate system changes its position over time, determining its orientation becomes nontrivial. To explore the physical meaning and possible experimental interpretation of this rotation, we illustrate the problem of defining the orientation of a moving frame using a simple example. The Wigner rotation, when interpreted in the context of the original inertial frame, is intimately related to length contraction  [^31] The length of a moving object depends not only on the structure of space-time but also on our choice of synchronization convention. As with the orientation of radiation phase fronts, such quantities do not possess an exact objective meaning. Under the standard clock synchronization, they are typically regarded as ”relativistic kinematic effects.” The Wigner rotation similarly arises as a coordinate effect, lacking an absolute, observer-independent significance..
@@
Suppose that, after performing a Lorentz boost along the $z$-axis with velocity $v_{z}$, an observer in the lab frame rotates the coordinate system by an angle $\vec{\delta\theta}=\vec{v_{z}}\times d\vec{v_{x}}/v_{z}^{2}$, such that the new $x_{r}$-axis is orthogonal to the vector $\vec{v_{z}}+d\vec{v_{x}}$, and the $z_{r}$-axis is aligned with it. Now, consider a rod oriented along the $x$-axis in the comoving frame. The motion takes place in the $(x,z)$-plane, and the rod is initially perpendicular to the velocity $\vec{v_{z}}$. After rotating the lab frame axes, the projection of the rod onto the $z_{r}$-axis becomes simply $l\delta\theta$, where $l$ is the length of the rod in the comoving frame, and also in the lab frame following the first boost along the $z$-axis. After a second infinitesimal Lorentz boost with velocity $d\vec{v_{x}}$, this projection undergoes Lorentz contraction and becomes $l\delta\theta/\gamma$ (Fig. 37). Assume now that the lab observer fixes the orientation of the comoving frame axes. In the ultrarelativistic limit $\gamma\to\infty$, the comoving frame axes become parallel to the rotated lab frame axes $(x_{r},z_{r})$, and thus the projection of the rod onto the $z_{r}$-axis vanishes. However, for an arbitrary (finite) velocity, the comoving frame axes are not aligned with the rotated lab frame axes. Due to the contracted projection, the angular deviation is $-\vec{\delta\theta}/\gamma$. Consequently, one can directly verify that the comoving frame axes are rotated relative to the original lab frame axes $(x,z)$ by an angle $\delta\theta-\delta\theta/\gamma$, which is precisely the Wigner rotation angle, as given by Eq.(20). This example illustrates how the Wigner rotation arises naturally from the composition of Lorentz transformations.
@@ 8.4 Discussion
The conventional approach to explaining spin dynamics in the lab frame is quite unusual. As previously noted, the laws of physics in the lab frame account for all physical phenomena, including observations made by non-inertial observers. In particular, spin orientation measurements with respect to the lab frame axes can be determined within the proper frame. From the lab frame perspective, the interpretation of spin rotation experiments is straightforward and can be derived from the predictions of the proper observer regarding the measurements made by the lab observer. Thus, the spin rotation measurements by a polarimeter in the lab frame is interpreted from the viewpoint of the proper observer as viewing of the lab observer. This inherent complexity has led textbook authors to derive incorrect expressions for Wigner rotation—an issue that will be discussed further in Chapter 17.
@@ 9.1 Explanation of Electron Aberration Based on Electrodynamics
To proceed, we transform the coordinates $(t,x,y,z)$ of the inertial observer $S$, moving with velocity $-v$ relative to the accelerated observer $S_{n}$, using the Galilean transformation: $x_{n}=x-vt$, $t_{n}=t$. Substituting these into the Minkowski metric, $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$, we obtain the so-called Langevin metric in the accelerated frame Eq. (15): $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$. This transformed metric allows us to study Maxwell’s equations in a non-inertial frame using the Galilean transformation with velocity $-v$. In the specific case where $\vec{B}=0$, the transformed fields to first order in $v/c$ are: $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$. Here, $\vec{v}$ is the velocity of the source in the original inertial frame. This result captures the leading-order relativistic correction to the magnetic field observed in the accelerated frame.
@@
[…]
@@ 9.2 Magnetic Field Measurements in an Accelerated Electron Source
We derived the electrodynamics equations in an accelerated frame by applying the inverse Galilean transformation (with velocity $-v$) of Maxwell’s equations. Under a Galilean transformation, and assuming $\vec{B}=0$ in the original frame, the transformed fields become $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$. The electric field $\vec{E}_{n}$ and the induced magnetic field $\vec{B}_{n}$ are both static, meaning they do not vary with time—the system is entirely time-independent. Many physicists familiar with standard electrodynamics might argue, based on textbook knowledge, that a static magnetic field must arise from electric currents, which in turn require moving charges. Since no charges are in motion in this case, the presence of a magnetic field appears paradoxical. This represents a fundamentally new kind of situation, distinct from traditional electrodynamics in an initial inertial frame.
How can that be? Even when the electron source and the observer are at rest in the accelerated frame, the time derivative $\partial/\partial t_{n}$ is generally nonzero. This can be readily demonstrated by transforming from the coordinates $(t,x,y,z)$ of an inertial observer $S$—moving with velocity $-v$ relative to the non-inertial observer $S_{n}$—to the coordinates $(t_{n},x_{n},y_{n},z_{n})$ of $S_{n}$. Using a Galilean transformation, the inertial coordinates can be related to the non-inertial ones as follows: $x_{n}=x-vt,~{}y_{n}=y,~{}z_{n}=z,~{}t_{n}=t$. This allows all fields to be re-expressed in terms of the variables $x_{n},y_{n},z_{n},t_{n}$. Taking partial derivatives accordingly, we find: $\partial/{\partial t_{n}}=\partial/{\partial t}+v\partial/{\partial x}$, $\partial/{\partial x_{n}}=\partial/{\partial x}$. Therefore, if the fields are time-independent in the inertial frame (i.e., $\partial/\partial t=0$), it follows that: $\partial/\partial t_{n}=v\partial/\partial x_{n}$. Thus, $\partial/\partial t_{n}$ is not necessarily zero, even though the source and observer are at rest in the non-inertial frame.
@@
To understand magnetic field measurements in an accelerated electron source, we begin by examining how a classical magnetometer behaves when observed from an accelerated frame. Since a full treatment is complex, we consider a simplified, conventional type of magnetometer. One practical approach to observing nuclear magnetism is through the phenomenon of nuclear magnetic resonance (NMR), where a proton resonance apparatus can serve effectively as a proton resonance magnetometer. We present a general analysis of this classical type of magnetometer, independent of its experimental precision. In the inertial lab frame, a current loop possessing a magnetic moment $\vec{m}$ acquires an electric dipole moment when it moves with velocity $\vec{v}$ relative to the lab. This dipole moment is given by: $\vec{d}=\vec{v}\times\vec{m}/c$. Since the magnetic properties of materials arise from atomic-scale current loops, a magnetized body undergoing acceleration relative to the lab frame should exhibit an induced electric polarization in that frame. As a result, the potential energy of a moving particle with intrinsic magnetic moment $\vec{m}$, in the presence of a magnetic field $\vec{B}$ and electric field $\vec{E}$, is given by: $U_{p}=-\vec{m}\cdot\vec{B}-\vec{d}\cdot\vec{E}$. To analyze this in an accelerated frame, we apply the Galilean transformation (with velocity $-\vec{v}$) to Maxwell’s equations. Under this transformation, assuming an initial zero electric dipole moment ($\vec{d}=0$), the magnetic and electric dipole moments in the accelerated frame become: $\vec{m}_{n}=\vec{m}$, $\vec{d}_{n}=-\vec{v}\times\vec{m}/c$. From vector algebra, we use the identity $(\vec{a}\times\vec{b})\cdot\vec{c}=\vec{a}\cdot(\vec{b}\times\vec{c})$ to evaluate terms in the potential energy expression: $(-\vec{v}\times\vec{E})\cdot\vec{m}=-\vec{v}\cdot(\vec{E}\times\vec{m})$, $(-\vec{v}\times\vec{m})\cdot\vec{E}=-\vec{v}\cdot(\vec{m}\times\vec{E})$. Thus, the additional contribution to the potential energy evaluates to zero, as expected. Consequently, a conventional magnetometer cannot detect the magnetic field $\vec{B}_{n}$ in the accelerated frame.
@@ 9.3 Aberration of Particles from a Moving Source
[…]
@@
Let us now return to the magnetic field measurements in the accelerated electron source, as shown in Fig. 39. In the lab frame, the electron source is in motion and its evolution is analyzed from the perspective of an inertial reference system. A key feature of this situation is that the velocity of the source carries real physical significance. Indeed, the problem is explicitly time-dependent. For a moving electron source, special relativity tells us that a magnetic field arises according to $\vec{B}=\vec{v}\times\vec{E}/c$, leading to a change in the transverse momentum of test particles. This change can be quantified as $\theta_{a}=p_{\perp}/p=v/V$, a result that remains valid even in the non-relativistic limit as $V^{2}/c^{2}\to 0$. This brings us to an interesting question: what does an inertial observer measure when using a conventional magnetometer? Given that charges in the source are moving, the resulting current produces a magnetic field described by $\vec{B}=\vec{v}\times\vec{E}/c$. This field is indeed detectable with a standard magnetometer. To understand this more concretely, consider the operation of an NMR magnetometer. When at rest in the lab frame, it measures the additional potential energy experienced by a proton with magnetic moment $\vec{m}$ in the field $\vec{B}$, given by $-\vec{m}\cdot\vec{B}\neq 0$. We come to the conclusion that the magnetic field generated by a moving source in an inertial frame is a real field in the sense we have defined it.
@@ 9.4 Inertial Frame View of Observations by a Non-Inertial Observer
The Wigner rotation is a relativistic kinematic effect in which the coordinate axes of a particle’s proper reference frame appear to rotate relative to an initial (Lorentz) inertial frame. This rotation arises due to successive Lorentz boosts in different directions. The Wigner rotation angle in the inertial frame can be expressed in vector form as $\vec{\theta_{w}}=(1-1/\gamma)\vec{V}\times\vec{v}/V^{2}$, where $\vec{V}$ is the vector of particle velocity in the initial Lorentz frame before acceleration, $\vec{v}$ is the vector of small velocity change ($v\ll V$) due to acceleration, $\gamma$ is the Lorentz factor associated with $V$, and $\theta_{w}$ is the Wigner rotation angle of the proper frame axes as observed from the (Lorentz) inertial frame. We now examine how this formula applies in the specific case where both the observer and the particle source are accelerated to a velocity $v$ in the inertial frame.
@@
Consider an electron gun–detector setup initially at rest in an inertial frame of reference. Suppose this system is then uniformly accelerated along the $x$-axis until it attains a velocity $v$. Now, from the perspective of an observer who remains at rest in the original inertial frame, the direction of the electron beam’s propagation appears to shift. This observer measures an angular displacement of the beam given by $\theta_{a}=v/V$, where $V$ is the velocity of the particles emitted by the gun. This effect, known as particle aberration, is a well-established phenomenon. A key distinction between relativistic and Newtonian kinematics emerges when we examine the behavior of the emitter–detector system under acceleration. From the inertial frame’s viewpoint, the accelerated setup undergoes a shearing motion: the detector no longer remains aligned with the electron beam after acceleration. For instance, if the gun–detector assembly was originally aligned along the $z_{n}$-axis, the detector fails to intercept the beam once the acceleration ends. In the accelerated frame, this misalignment is reflected in the angular deviation, and we obtain an expression for the aberration increment: $\theta_{a}=-\theta_{w}=-(1-1/\gamma)v/V$. In the special case of light, where $V\to c$, this expression simplifies to: $\theta_{a}=-v/c$.

### v12

@@ 8.3 Measuring Wigner rotation
Here $d\vec{v}$ represents the infinitesimal change in velocity due to acceleration, $d\Phi$ is the Wigner rotation angle describing the rotation of the particle’s proper frame relative to the Lorentz lab frame, and $d\theta$ is the infinitesimal orbital angle. A natural question arises: how can this rotation be observed or measured? Since a moving coordinate system changes its position over time, determining its orientation becomes nontrivial. To explore the physical meaning and possible experimental interpretation of this rotation, we illustrate the problem of defining the orientation of a moving frame using a simple example. The Wigner rotation, when interpreted in the context of the original inertial frame, is intimately related to length contraction.  [^31] The length of a moving object depends not only on the structure of space-time but also on our choice of synchronization convention. As with the orientation of radiation phase fronts, such quantities do not possess an exact objective meaning. Under the standard clock synchronization, they are typically regarded as ”relativistic kinematic effects.” The Wigner rotation similarly arises as a coordinate effect, lacking an absolute, observer-independent significance.
@@
Suppose that, after performing a Lorentz boost along the $z$-axis with velocity $v_{z}$, an observer in the lab frame rotates the coordinate system by an angle $\vec{\delta\theta}=\vec{v_{z}}\times d\vec{v_{x}}/v_{z}^{2}$, such that the new $x_{r}$-axis is orthogonal to the vector $\vec{v_{z}}+d\vec{v_{x}}$, and the $z_{r}$-axis is aligned with it. Now, consider a rod oriented along the $x$-axis in the comoving frame. The motion takes place in the $(x,z)$-plane, and the rod is initially perpendicular to the velocity $\vec{v_{z}}$. After rotating the lab frame axes, the projection of the rod onto the $z_{r}$-axis becomes simply $l\delta\theta$, where $l$ is the length of the rod in the comoving frame, and also in the lab frame following the first boost along the $z$-axis. After a second infinitesimal Lorentz boost with velocity $d\vec{v_{x}}$, this projection undergoes Lorentz contraction and becomes $l\delta\theta/\gamma$ (Fig. 37). Assume now that the lab observer fixes the orientation of the comoving frame axes. In the ultrarelativistic limit $\gamma\to\infty$, the comoving frame axes become parallel to the rotated lab frame axes $(x_{r},z_{r})$, and thus the projection of the rod onto the $z_{r}$-axis vanishes. However, for an arbitrary (finite) velocity, the comoving frame axes are not aligned with the rotated lab frame axes. Due to the contracted projection, the angular deviation is $\vec{\delta\theta}/\gamma$. Consequently, one can directly verify that the comoving frame axes are rotated relative to the original lab frame axes $(x,z)$ by an angle $\delta\theta-\delta\theta/\gamma$, which is precisely the Wigner rotation angle, as given by Eq.(20). This example illustrates how the Wigner rotation arises naturally from the composition of Lorentz transformations.
@@ 8.4 Discussion
The conventional approach to explaining spin dynamics in the lab frame is quite unusual. As previously noted, the laws of physics in the lab frame account for all physical phenomena, including observations made by non-inertial observers. In particular, spin orientation measurements with respect to the lab frame axes can be determined within the proper frame. From the lab frame perspective, the interpretation of spin rotation experiments is straightforward and can be derived from the predictions of the proper observer regarding the measurements made by the lab observer. Thus, the spin rotation measurements by a polarimeter in the lab frame is interpreted from the viewpoint of the proper observer as viewing of the lab observer. This inherent complexity has led textbook authors to derive incorrect expressions for Wigner rotation—an issue that will be discussed further in Chapter 18.
@@ 9.1 Explanation of Electron Aberration Based on Electrodynamics
To proceed, we transform the coordinates $(t,x,y,z)$ of the inertial observer $S$, moving with velocity v-v relative to the accelerated observer $S_{n}$, using the Galilean transformation: $x_{n}=x-vt$, $t_{n}=t$. Substituting these into the Minkowski metric, $ds^{2}=c^{2}dt^{2}-dx^{2}-dy^{2}-dz^{2}$, we obtain the so-called Langevin metric in the accelerated frame Eq. (15): $ds^{2}=c^{2}(1-v^{2}/c^{2})dt_{n}^{2}-2vdx_{n}dt_{n}-dx_{n}^{2}-dy_{n}^{2}-dz_{n}^{2}$. This transformed metric allows us to study electrodynamics equations in a non-inertial frame using the Galilean transformation with velocity v-v. In the specific case where $\vec{B}=0$, the transformed fields to first order in $v/c$ are: $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$. Here, $\vec{v}$ is the velocity of the source in the original inertial frame. This result captures the leading-order relativistic correction to the magnetic field observed in the accelerated frame.
@@
The accelerated observer experimentally determines the Lorentz force $\vec{F}_{n}$ acting on a charge $e$ moving with velocity $\vec{V}$ in the region of the electron gun, where electric and magnetic fields $\vec{E}_{n}$ and $\vec{B}_{n}$ are present. The force is given by: $\vec{F}_{n}=e\vec{E}-e\vec{V}\times\vec{v}\times\vec{E}/c^{2}$. In the special case where $\vec{v}\cdot\vec{V}=0$ and $\vec{E}\cdot\vec{V}=EV$, the magnetic term simplifies to: $e\vec{V}\times\vec{v}\times\vec{E}/c^{2}=-\vec{v}(e\vec{V}\cdot\vec{E}/c^{2})$. Using the relativistically correct equation of motion, $\vec{F}_{n}=d\vec{p}/dt$, we expect that after a short time interval $\d t$, the emitted electron will acquire a transverse momentum in the accelerated frame given by: $\d\vec{p}_{\perp}=-\vec{v}(e\vec{V}\cdot\vec{E}/c^{2})\d t$. The differential change in kinetic energy of the accelerated electrons is: $\d T=\vec{F}\cdot\vec{V}\d t=\vec{F}\cdot\d\vec{s}$. Upon integration, this yields: $T=e\int\vec{E}\cdot\d\vec{s}=mc^{2}/\sqrt{1-V^{2}/c^{2}}-mc^{2}$. Given our condition $\vec{v}\cdot\vec{V}=0$, the aberration angle increment becomes: $\theta_{a}=p_{\perp}/p$, where the total momentum of the accelerated electron is: $p=mV/\sqrt{1-V^{2}/c^{2}}$. Substituting, we find: $\theta_{a}=-(1-1/\gamma)v/V$. Expressed in vector form, the aberration increment is: $\vec{\theta}_{a}=-(1-1/\gamma)\vec{V}\times\vec{v}/V^{2}$.
@@ 9.2 Magnetic Field Measurements in an Accelerated Electron Source
We derived the electrodynamics equations in an accelerated frame by applying the inverse Galilean transformation (with velocity v-v) of Maxwell’s equations. Under a Galilean transformation, and assuming $\vec{B}=0$ in the original frame, the transformed fields become $\vec{E}_{n}=\vec{E}$, $\vec{B}_{n}=-\vec{v}\times\vec{E}/c$. The electric field $\vec{E}_{n}$ and the induced magnetic field $\vec{B}_{n}$ are both static, meaning they do not vary with time—the system is entirely time-independent. Many physicists familiar with standard electrodynamics might argue, based on textbook knowledge, that a static magnetic field must arise from electric currents, which in turn require moving charges. Since no charges are in motion in this case, the presence of a magnetic field appears paradoxical. This represents a fundamentally new kind of situation, distinct from traditional electrodynamics in an initial inertial frame.
How can that be? Even when the electron source and the observer are at rest in the accelerated frame, the time derivative $\partial/\partial t_{n}$ is generally nonzero. This can be readily demonstrated by transforming from the coordinates $(t,x,y,z)$ of an inertial observer SS—moving with velocity v-v relative to the non-inertial observer $S_{n}$—to the coordinates $(t_{n},x_{n},y_{n},z_{n})$ of $S_{n}$. Using a Galilean transformation, the inertial coordinates can be related to the non-inertial ones as follows: $x_{n}=x-vt,~y_{n}=y,~z_{n}=z,~t_{n}=t$. This allows all fields to be re-expressed in terms of the variables $x_{n},y_{n},z_{n},t_{n}$. Taking partial derivatives accordingly, we find: $\partial/{\partial t_{n}}=\partial/{\partial t}+v\partial/{\partial x}$, $\partial/{\partial x_{n}}=\partial/{\partial x}$. Therefore, if the fields are time-independent in the inertial frame (i.e., $\partial/\partial t=0$), it follows that: $\partial/\partial t_{n}=v\partial/\partial x_{n}$. Thus, $\partial/\partial t_{n}$ is not necessarily zero, even though the source and observer are at rest in the non-inertial frame.
@@
To understand magnetic field measurements in an accelerated electron source, we begin by examining how a classical magnetometer behaves when observed from an accelerated frame. Since a full treatment is complex, we consider a simplified, conventional type of magnetometer. One practical approach to observing nuclear magnetism is through the phenomenon of nuclear magnetic resonance (NMR), where a proton resonance apparatus can serve effectively as a proton resonance magnetometer. We present a general analysis of this classical type of magnetometer, independent of its experimental precision. In the inertial lab frame, a current loop possessing a magnetic moment $\vec{m}$ acquires an electric dipole moment when it moves with velocity $\vec{v}$ relative to the lab. This dipole moment is given by: $\vec{d}=\vec{v}\times\vec{m}/c$. Since the magnetic properties of materials arise from atomic-scale current loops, a magnetized body undergoing acceleration relative to the lab frame should exhibit an induced electric polarization in that frame. As a result, the potential energy of a moving particle with intrinsic magnetic moment $\vec{m}$, in the presence of a magnetic field $\vec{B}$ and electric field $\vec{E}$, is given by: $U_{p}=-\vec{m}\cdot\vec{B}-\vec{d}\cdot\vec{E}$. To analyze this in an accelerated frame, we apply the Galilean transformation (with velocity $\vec{v}$) to Maxwell’s equations. Under this transformation, assuming an initial zero electric dipole moment ($\vec{d}=0$), the magnetic and electric dipole moments in the accelerated frame become: $\vec{m}_{n}=\vec{m}$, $\vec{d}_{n}=-\vec{v}\times\vec{m}/c$. From vector algebra, we use the identity $(\vec{a}\times\vec{b})\cdot\vec{c}=\vec{a}\cdot(\vec{b}\times\vec{c})$ to evaluate terms in the potential energy expression: $(-\vec{v}\times\vec{E})\cdot\vec{m}=-\vec{v}\cdot(\vec{E}\times\vec{m})$, $(-\vec{v}\times\vec{m})\cdot\vec{E}=-\vec{v}\cdot(\vec{m}\times\vec{E})$. Thus, the additional contribution to the potential energy evaluates to zero, as expected. Consequently, a conventional magnetometer cannot detect the magnetic field $\vec{B}_{n}$ in the accelerated frame.
@@ 9.3 Aberration of Particles from a Moving Source
An inertial observer experimentally measures the Lorentz force $\vec{F}$ acting on a charge $e$ moving with velocity $\vec{V}$ in the region of an electron gun, where both electric and magnetic fields, $\vec{E}$ and $\vec{B}$, are present. The force is given by: $\vec{F}=e\vec{E^{\prime}}+e\vec{V}\times\vec{v}\times\vec{E^{\prime}}/c^{2}$, where $\vec{E}^{\prime}$ is the electric field in the coordinate system moving with velocity $\vec{v}$ relative to the inertial frame. In the special case where $\vec{v}\cdot\vec{V}=0$ and $\vec{E}^{\prime}\cdot\vec{V}=E^{\prime}V$, the magnetic component of the force simplifies to: $e\vec{V}\times\vec{v}\times\vec{E^{\prime}}/c^{2}=\vec{v}(e\vec{V}\cdot\vec{E^{\prime}}/c^{2})$. The differential change in the kinetic energy of the accelerated electrons is: $dT=\vec{F}\cdot\vec{V}dt=\vec{F}\cdot d\vec{s}$. Integrating this expression yields the total kinetic energy: $T=e\int\vec{E^{\prime}}\cdot d\vec{s}=mc^{2}/\sqrt{1-V^{2}/c^{2}}-mc^{2}$. The total transverse momentum $p_{\perp}$ of the accelerated electron beam, to first order in $v/c$, is given by: $p_{\perp}=p_{0}+\Delta p_{\perp}=mv+vT/c^{2}=mv/\sqrt{1-V^{2}/c^{2}}$, where $p_{0}=mv$ is the initial transverse momentum of the emitted electron. Since we have chosen $\vec{v}\cdot\vec{V}=0$, the aberration angle $\theta_{a}$ becomes: $\theta_{a}=p_{\perp}/p$, where $p=mV/\sqrt{1-V^{2}/c^{2}}$ is the momentum of the accelerated electron. Substituting, we obtain: $\theta_{a}=v/V$. This can be written in vector form as: $\vec{\theta}_{a}=\vec{V}\times\vec{v}/V^{2}$. It is important to note that this result holds for an arbitrary Lorentz factor $\gamma=1/\sqrt{1-V^{2}/c^{2}}$.
@@
Let us now return to the magnetic field measurements in the accelerated electron source, as shown in Fig. 39. In the lab frame, the electron source is in motion and its evolution is analyzed from the perspective of an inertial reference system. A key feature of this situation is that the velocity of the source carries real physical significance. Indeed, the problem is explicitly time-dependent. For a moving electron source, special relativity tells us that a magnetic field arises according to $\vec{B}=\vec{v}\times\vec{E}/c$, leading to a change in the transverse momentum of test particles. This change can be quantified as $\theta_{a}=p_{\perp}/p=v/V$, a result that remains valid even in the non-relativistic limit as $V^{2}/c^{2}\to 0$. This brings us to an interesting question: what does an inertial observer measure when using a conventional magnetometer? Given that charges in the source are moving, the resulting current produces a magnetic field described by $\vec{B}=\vec{v}\times\vec{E}/c$. This field is indeed detectable with a standard magnetometer. To understand this more concretely, consider the operation of an NMR magnetometer. When at rest in the lab frame, it measures the additional potential energy experienced by a proton with magnetic moment $\vec{m}$ in the field $\vec{B}$, given by $\vec{m}\cdot\vec{B}\neq 0$. We come to the conclusion that the magnetic field generated by a moving source in an inertial frame is a real field in the sense we have defined it.
@@ 9.4 Inertial Frame View of Observations by a Non-Inertial Observer
The Wigner rotation is a relativistic kinematic effect in which the coordinate axes of a particle’s proper reference frame appear to rotate relative to an initial (Lorentz) inertial frame. This rotation arises due to successive Lorentz boosts in different directions. The Wigner rotation angle in the inertial frame can be expressed in vector form as $\vec{\theta_{w}}=(1-1/\gamma)\vec{V}\times\vec{v}/V^{2}$, where $\vec{V}$ is the vector of particle velocity in the initial Lorentz frame before acceleration, $\vec{v}$ is the vector of small velocity change ($v\ll V$) due to acceleration, $\gamma$ is the Lorentz factor associated with $V$, and $\theta_{w}$ is the Wigner rotation angle of the proper frame axes as observed from the (Lorentz) inertial frame.  [^32] This expression for the Wigner rotation is derived in Chapter 8 for the limiting case of an infinitesimal rotation angle, $\delta\theta=v/V$, corresponding to a small change in the particle’s velocity vector. The derivation is based on the composition of Lorentz boosts in the form $L=L(\vec{v})L(\vec{V})$. In the context of an accelerated electron source, the relevant velocity composition arises from $L=L(\vec{V})L(\vec{v})$. A careful analysis shows that both orderings yield the same Wigner rotation expression, due to the infinitesimal nature of the boost $L(\vec{v})$. We now examine how this formula applies in the specific case where both the observer and the particle source are accelerated to a velocity $v$ in the inertial frame.
@@
Consider an electron gun–detector setup initially at rest in an inertial frame of reference. Suppose this system is then uniformly accelerated along the $x$-axis until it attains a velocity $v$. Now, from the perspective of an observer who remains at rest in the original inertial frame, the direction of the electron beam’s propagation appears to shift. This observer measures an angular displacement of the beam given by $\theta_{a}=v/V$, where $V$ is the velocity of the particles emitted by the gun. This effect, known as particle aberration, is a well-established phenomenon. A key distinction between relativistic and Newtonian kinematics emerges when we examine the behavior of the emitter–detector system under acceleration. From the inertial frame’s viewpoint, the accelerated setup undergoes a shearing motion: the detector no longer remains aligned with the electron beam after acceleration. For instance, if the gun–detector assembly was originally aligned along the $z_{n}$-axis, the detector fails to intercept the beam once the acceleration ends. In the accelerated frame, this misalignment is reflected in the angular deviation, and we obtain an expression for the aberration increment: $\theta_{a}=-\theta_{w}=-(1-1/\gamma)v/V$. In the special case of light this expression simplifies to: $\theta_{a}=-v/c$.  [^33] It’s straightforward to verify that the infinitesimal Wigner rotation angle $\theta_{w}=(1-1/\gamma)v/V$ is always less than $v/c$ regardless of the particle velocity $V$. Specifically, $\theta_{w}=(v/c)(1-\sqrt{1-\beta^{2}})/\beta$, where $\beta=V/c$ and function $(1-\sqrt{1-\beta^{2}})/\beta$ is strictly less than 1 for $0<\beta<1$. As discussed in Chapter 5, there is an inherent uncertainty in the relative position along the $z_{n}$-axis in the $x$-direction, amounting to $z_{n}v/c$, due to the ambiguity in the timing of acceleration. This uncertainty arises because the relevant events are space-like separated. Consequently, the orientation of the $z_{n}$ axis relative to the inertial frame’s $z$-axis carries an angular uncertainty of $v/c$. The Wigner inclination angle $\theta_{w}$ always lies within this uncertainty bound.